OpenStudy (anonymous):

I'm feeling really frustrated that I can't remember this.... but how can I get this into a form where I can use quadratic formula??

OpenStudy (anonymous):

I have \[\frac{x ^{2} }{ 0.0749 - x } =1.80\]

OpenStudy (anonymous):

I need to change it to where it it in the form of Ax^2 +Bx + C = 0